A 68.5-kg skater moving initially at 2.40 m/s on rough horizontal ice comes to rest uniformly in 3.52 s due to friction from the ice. What force does friction exert on the skater?


1
Expert's answer
2021-12-13T16:55:03-0500

Solution:


To determine the friction force exerted in the skater, we use Newton's Second law of motion which relates force, acceleration, and mass. It is expressed as follows: 


F = m (a) 


where

m = mass of the body = 68.5 kg 


The acceleration, a, of the moving body is calculated by dividing the initial velocity by the total time. Thus,


"a=\frac{2.40(\frac{m}{s})}{3.52 s}=0.682(\frac{m}{s^2})"


Substituting the values to the equation of force, we will have


F = m (a) = 68.5 kg ( 0.682 "(\frac{m}{s^2})" ) = 46.7 N 


Thus, the friction force that is being exerted by the ice to the skater ould be 46.7 N.